OpenStudy (anonymous):

Help solving this problem. (3/c^2-4c)-(9/2c^2+3c)=(2/2c^2-5c-12)

OpenStudy (anonymous):

\[\frac{ 3 }{ c ^{2} -4c}-\frac{ 9 }{ 2c ^{2}+3c }= \frac{ 2 }{ 2c^{2}-5c-12 }\]